OpenStudy (anonymous):

Given the function g(x) = 8x - 2, find g(-2)

OpenStudy (anonymous):

i think its 8?

Parth (parthkohli):

i tackled a similar question before. you just have to substitute.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 8(-2) - 2 }\) \( \color{Black}{\Rightarrow -16 - 2 }\) \( \color{Black}{\Rightarrow -18 }\)

OpenStudy (anonymous):

oh okay i did it wrong thanks

Parth (parthkohli):

well no worries I do it wrong too lol

OpenStudy (anonymous):

?

OpenStudy (australopithecus):

g(x) = 8x - 2, find g(-2) g(-2) sub -2 into x g(-1) sub -1 into x etc so your equation is g(-2) = 8(-2) - 2 g(-1) = 8(-1) - 2 g(0) = 8(0) - 2 etc...

OpenStudy (australopithecus):

the reason we use this notation is because it would be confusing to work with multiple functions if we made them all equal to y otherwise thus we call the function a letter like g or f. Also, we write the variable that y changes in respect to next to the name of the function so that we can talk about a specific point in the function in an easier fashion, it also tells us the variable y is changing in respect to in a function. that is why we use this notation g(x), f(x)...